1. Which of the following is NOT a component of a vulnerability chart?

a.

​concealment possibilities

b.

​identification of the suspected perpetrator

c.

​listing assets that are missing

d.

​internal controls that had to be compromised

2.

Which of the following is a benefit of creating a vulnerability chart?

a.

​It shows management where controls are weak.

b.

​It lists/coordinates the steps the investigator must take to uncover the perpetrator.

c.

​It forces the investigators to explicitly consider all aspects of a fraud act.

d.

​It conveys a warning to all employees that fraud will not be tolerated.

3.

Which of the following is a limitation to the Net Worth Method?

a.

​The Net Worth Method only gives an estimate of Net Worth increases or decreases.

b.

​It only includes assets and liabilities that can be discovered.

c.

​The resulting amounts presented in court are usually lower than the actual theft amount.

d.

​When presented with the Net Worth amount(s), suspects sometimes confess.

4.

Data Set 9-1

Year 1 Year 2 Year 3
Home

220000

220000

220000

Car

25000

25000

50000

Investments

40000

50000

65000

Bank account

3000

7000

10000

Total Assets

288000

302000

345000

Home loan

200000

150000

100000

Auto loan

20000

0

0

Student loan

60000

20000

0

Total Liabilities

280000

170000

100000

Salary

95000

105000

Investments

4000

5000

Inheritance

6000

0

Total Income

105000

110000

Home loan

13200

13200

Auto loan

9600

9600

Other living

20000

22000

Wedding

4000

0

Total Expenses

46800

44800

Refer Data Set 9-1. Calculate the net worth of the individual at the end of year 1.

a.

$0

b.

$288,000

c.

$3,000

d.

$8,000

5.

Year 1 Year 2 Year 3
Home

220000

220000

220000

Car

25000

25000

50000

Investments

40000

50000

65000

Bank account

3000

7000

10000

Total Assets

288000

302000

345000

Home loan

200000

150000

100000

Auto loan

20000

0

0

Student loan

60000

20000

0

Total Liabilities

280000

170000

100000

Salary

95000

105000

Investments

4000

5000

Inheritance

6000

0

Total Income

105000

110000

Home loan

13200

13200

Auto loan

9600

9600

Other living

20000

22000

Wedding

4000

0

Total Expenses

46800

44800

Refer Data Set 9-1. The income from unknown sources for years 2 and 3 respectively are:

a.

$255,200; $300,200

b.

$171,200; $307,200

c.

$73,800; $179,800

d.

$65,800; $47,800
Answer #1

1. option C

​listing assets that are missing (all other given components are related to vulnerability chart)

2. option C.

It forces the investigators to explicitly consider all aspects of a fraud act. (vulnerability chart is created for internal control and thus it becomes necessary for investigators to consider all factors possible for fraud)

3. option B

​It only includes assets and liabilities that can be discovered. (new worth = assets –liabilites)

4. option D

$8000

Net worth at the end of Year 1 = 288,000 - 280,000 = $8,000

5. option D

$65,800; $47,800

Funds from unknown sources = Assets - liabilities - prior year's net worth + living expenses - funds from known sources.

Year 2: 302,000 - 170,000 - (288,000 - 280,000) + 46,800 - 105,000 = 65,800

Year 3: 345000-100000 –(302000-170000)+44800-110000 = 47800